Solve the pair of equation 2y^-1 -3z^-1=4 and 4y^-1+z^-1=1

1
Expert's answer
2022-01-31T16:02:35-0500

Let us change the variables: x = 1/y, t = 1/z. Therefore, our equations take form

"\begin{cases}
2x -3t = 4, \\ 4x + t = 1. 
\end{cases}"

From the second equation we obtain "t = 1-4x." Next, we substitute it in the first equation and obtain

"2x - 3(1-4x) = 4, \\
14x = 7, \\
x = 0.5"

Therefore, "t = 1-4\cdot 0.5 = -1."


"y = 1/x = 1/0.5 = 2, \\
z = 1/t = 1/(-1) = -1."